Consider the vectors = (1/72,1172) and J= (-1112,112). Write the vector (2, - 4) in terms of I and J. (2. - 4) = O + ( DJ
Answer #1

Les [2, (2,-4)= l it my .:(2,-4) 20 sta to 3 pm L - a ta) e the ☺ < e-m V2 v2 Comparing wey et بدا مع l-bu also, -4 = r2 V2 esolving e-m lem - 252 -ur2 21 = -282 rea-v2 al som =-482- = -4/2+r2 2-32 m=-382 Th (1-352 ::22,-4)= (-12) Pt (Auswem